Kose, Inc., has a target debt-equity ratio of .57. Its WACC is 10.4 percent, and the tax rate is 25 percent.

a.

If the company’s cost of equity is 15 percent, what is its pretax cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. If instead you know that the aftertax cost of debt is 4.6 percent, what is the cost of equity? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a)

Weight of equity = 1 / (1 + 0.57) = 0.6369

Weight of debt = 1 - 0.6369 = 0.3631

WACC = Weight of equity*cost of equity + weight of debt*after tax cost of debt

0.104 = 0.6369*0.15 + 0.3631*Pre tax cost of debt*(1 - 0.25)

0.104 = 0.095535 + 0.272325*Pre tax cost of debt

Pres tax cost of debt = 0.0311 or 3.11%

b)

WACC = Weight of equity*cost of equity + weight of debt*after tax cost of debt

0.104 = 0.6369*cost of equity + 0.3631*0.046

0.104 = 0.6369*cost of equity + 0.016703

cost of equity = 0.1371 or 13.71%